Please, need help again.
Solve the following equation for w
2 (w - 3) + 2w = 14
Simply ou answer as much as possible

Solving a linear equation with several occurances of the variable
Solve for x
7/4x + 4/3 = 1/4x + 7

Solving a linear equation with several occurance of the variable
Solve the following equation for x
-5 (x + 8) = 4x + 32

Solving a linear equation with several occurances of the variable
Solve the following equation for x
-2(x - 4) + 7 = -6(8x - 6) + x

Solving a linear equation with several occurances of the variable
Solve the following equation for t
8t - 4/5 + 9t - 2/2 = 47
Simply your answer as much as possible

Answer by zeynep(43)

2 (w - 3) + 2w = 14
Let's remove the parantheses by using distrubitive property;
2w-6+2w=14
4w-6=14 Add additive inverse of -6 to both sides;
4w-6+6=14+6
4w=20 Divide both sides by 4;
w=5
7/4x + 4/3 = 1/4x + 7
Let's get rid of those fractions by multiplyying both sides by the LCD of 12;
12.(7/4x + 4/3) =12.( 1/4x + 7)
21x+16=3x+84
21x+16-3x=3x+84-3x
18x+16=84
18x+16-16=84-16
18x=68 divide both sides by 18;
x=68/18 if you simplify 68/18;
x=34/9

-5 (x + 8) = 4x + 32
-5x-40=4x+32
-5x-40-4x=4x+32-4x
-9x-40=32
-9x-40+40=32+40
-9x=72 divide both sides by -9;
x=-8

-2(x - 4) + 7 = -6(8x - 6) + x
-2x+8+7=-48x+36+x
-2x+15=-47x+36
-2x+15+47x=-47x+36+47x
45x+15=36
45x+15-15=36-15
45x=21 divide both sides by 45;
x=21/45 if you simplify 21/45;
x=7/15
8t - 4/5 + 9t - 2/2 = 47
Do you mean;
(8t - 4)/5 + (9t - 2)/2 = 47 ?